• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name strauch

    Log entry time 07:59:12 on June30,2000

    Entry number 45275

    keyword=Owl Shift Summary




    Owl Shift Summary,

    00:00 continuation of data taking
    01:25 Data Ac problem: No count rate, normal dead time, beam / detectors fine.
    01:30 continuation of data taking
    06:30 moved E-arm to 18.070 deg
    06:45 PS1 & PS3 single data
    07:00 moved H-arm to 47.092 deg
    while changing the fields: HRSH Q2 & HRSE Q3 trip,
    went to controlled access, Ed Folts reseted magnets



    Run Summary

    #2662 : 1.3 M events
    #2663 : 1.3 M events
    #2664 : Coda problem
    #2665 : 1.3 M events
    #2666 : 1.3 M events
    #2667 : 1.3 M events
    #2668 : 1.3 M events
    #2669 : 1.3 M events
    #2670 : 1.3 M events
    #2671 : 1.3 M events
    #2672 : 1.3 M events
    #2673 : 1.3 M events
    #2674 : 1.3 M events
    #2675 : 0.6 M events

    sum(chrg_sum_us3) = 3.87 C


    optics and q-pointing sequence

    run events conditions
    ==== ===== ========
    2676 160k thetae=18.07,pe=3.66,ps1=23,ps3=80,others high; HRSh same
    2677 277k same (2676 interrupted by brief beam outage)